2. The pion has an average lifetime of 26.0ns when at rest. For it to travel 10.0m, how fast must it move?
Throughout this homework assignment, you may use e to denote the speed of light in both your work and final answers without using its numerical value. (This is actually preferred unless the problem requires the numerical value of e.)
Answer #1

answer) we have the formula

t=to\sqrt{}1-v2/c2=to/1-v2/c2

and we know speed=distance/time

v=l/t

v=l/t2+(l2/c2)

so plugging in the values

v=10/(26*10-9)2+(102/c2)=2.366*108m/s/3*108m/s=0.789c

so the answer is 0.789c